Using the formula y=20.2 log(x), where y=the height of the tree & x=time in years; predict how old the tree will be when it is 15m tall.
I want to change it into x=20.2^15, but the answer I get seems very extreme.

Answer by rothauserc(4718)

y = 20.2 log(x)
log(x) = y / 20.2
we are given y = 15
log(x) = 15 / 20.2
0.742574257 = log(x)
now use definition of logarithm
x = 10^(0.742574257)
x = 5.528079212
the tree is approx 5.5 years old